Consumer buys 10 units of Good A when the price of Good B is $5. When the price of Good B rises to $6 (the price of Good A remaining unchanged) the consumer buys 14 units of Good A.

Part A
Using an appropriate formula, calculate this Consumer’s cross Elasticity of demand for Good A. Show your working.
1
Expert's answer
2021-08-31T16:20:50-0400

we will find it by the formula:

"Eab=\frac{\frac{\Delta Qa}{Qa}}{\frac{\Delta Pb}{Pb}}=\frac{\frac{14-10}{10}}{\frac{6-5}{5}}=2"